OpenStudy (anonymous):

A camper has 3 jackets, 6 sleeping bags, and 14 different packages of food. An outfit consists of 2 jackets, 2 sleeping bags, and 3 packets of food. How many ways can the trip be outfitted?______

OpenStudy (anonymous):

much easier than the last one

OpenStudy (anonymous):

Similar to the other one right?

OpenStudy (anonymous):

\[\dbinom{3}{2}\times \dbinom{6}{2}\times \dbinom{14}{3}\]

OpenStudy (anonymous):

is it clear how to compute these numbers, because in fact all i really did was restate the question in notation

OpenStudy (anonymous):

how exactly do i solve that notation? multiply? Cross multiply?

OpenStudy (anonymous):

sorry it has been forever since i had a math class now im in this finite course:)

OpenStudy (anonymous):

maybe you have seen it this way instead \[3C2\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3x2x1 ----- 2x1 ?

OpenStudy (anonymous):

number of ways to choose 2 from a set of 3 is 3 number of ways to choose 2 from a set of 6 is \[\frac{6\times 5}{2}=15\]

OpenStudy (anonymous):

how do u get the 6 and 5

OpenStudy (anonymous):

yes that is one way to think about it. the other way is more basic. if you have three items and wish to select one, you pretty clearly have 3 choices. if you have three items and wish to select 2, you have the same three choice ( like selecting the one to exclude)

OpenStudy (anonymous):

you have 6 sleeping bags and wish to select 2 of them, six choices for first bag, 5 for second and you have to divide by two because order doesn't count

OpenStudy (anonymous):

the last one is therefore \[\frac{14\times 13\times 12}{3\times 2}\]

OpenStudy (anonymous):

so there are only two then correct?

OpenStudy (anonymous):

two what?

OpenStudy (anonymous):

are there two cases for this

OpenStudy (anonymous):

for this problem it is a straightforward multiplication \[\dbinom{3}{2}\times \dbinom{6}{2}\times \dbinom{14}{3}\] \[3\times 15\times 364\]

OpenStudy (anonymous):

wow you are a genius 16380

OpenStudy (anonymous):

but....im not sure how you got 3x15x364 what do i do with the numbers inside ( ) exactly?

OpenStudy (anonymous):

got disconnected